OpenStudy (anonymous):

calculate the pH of a 0.1 M solution of baking soda. showing step by step.

OpenStudy (aaronq):

Baking soda acts as a base at pH =7 so write an equation for the process and it's equilibrium expression. Make an ICE table and plug the values from E into the eq. expression. use \([OH^-]\) to find pH.

OpenStudy (anonymous):

how would i write the equation for the process and its equilibrium expression.

OpenStudy (aaronq):

\(HCO_3^-+H_2O\rightleftharpoons OH^-+H_2CO_3\) \(K=\dfrac{[products]}{[reactants]}=\dfrac{[OH^-][H_2CO_3]}{[HCO_3][H_2O]}\) but it reduces to: \(K=\dfrac{[OH^-][H_2CO_3]}{[HCO_3]}\)

OpenStudy (anonymous):

thank you so much that was lots of help.

OpenStudy (aaronq):

no problem, glad i could help

OpenStudy (anonymous):

:)